OpenStudy (anonymous):

helppp Peter has a briefcase that has a 3-digit combination lock where each digit can range from 0 - 9. For his birthday, Peter's wife, Sandy, decided to surprise him by placing Arizona Cardinal tickets in his briefcase for him to find once he was at work. However, Sandy does not know the combination to the briefcase. What is the probability that Sandy guesses the correct combination to the briefcase on her first attempt?

OpenStudy (anonymous):

You have 3-digits, and each digit can range to 0 to 9 so.. ALL POSSIBLE COMBINATIONS OF NUMBERS for the first digit you have 10 possibilities same for the second one and the third. NUMBER OF ATTEMPTS = 1 Probability = attempts / possibilities = 1/10.10.10 = 1/1000